OpenStudy (kainui):

\[\LARGE \frac{d^n}{dx^n}[e^{(e^{i \frac{2\pi}{n}})x}]=e^{(e^{i \frac{2\pi}{n}})x} \] Can we go deeper? e^e^e^(i...).... ?

OpenStudy (kainui):

\[\LARGE \frac{d^n}{dx^n}[e^{(e^{i \frac{2\pi}{n}})x}]=e^{(e^{i \frac{2\pi}{n}})x} \]